Main INDEX
Monthly INDEX
PREV
NEXT
User name LeRose
Log entry time 16:02:57 on April16,2008
Entry number 232442
keyword=Shift Summary
Day Shift 4/16/2008
J. LeRose SL
J. Udias TO
C. Fernandez-Ramirez 3rd
Kinematics F
LHRS 14.5 degrees P0 = 997 MeV/c
BigBite 390 Amps 48 degrees
beam_e0=1.193
08:27 MCC takes beam away for ~5 min. for a QE measurement
10:57 ROC5 crashed stopped runs restarted
11:01 back in business
12:30 J-P starts target temperature adjustments
12:52 J-P finishes his adjustments
14:05 ROC5 crashed stopped runs restarted
14:40-46 Doug and Vince do trigger test (see halog 232422)
Except for 6 minutes of ROC5 resetting almost all BANU is normal overhead
from starting and stopping more than 40 data runs runs (<1minute/run)
runs
07:54 3138 Production target_type=Loop 1 LH2 6 cm
07:57 20676 Production target_type=Loop 1 LH2 6 cm
08:07 3139 Production target_type=Loop 1 LH2 6 cm
08:20 3140 Production target_type=Loop 1 LH2 6 cm
08:39 3141 Production target_type=Loop 1 LH2 6 cm
08:51 3142 Production target_type=Loop 1 LH2 6 cm
09:01 3143 same
09:12 3144 same
09:25 3145 same
09:36 3146 same
09:46 3147 same
09:58 3148 same
10:09 3149 same
10:20 3150 same
10:31 3151 same
10:42 3152 same
10:54 3153 same
10:57 ROC5 crash
11:00 20677 Production target_type=Loop 1 LH2 6 cm
11:00 3154 Production target_type=Loop 1 LH2 6 cm
11:10 3155 same
11:20 3156 same
11:35 3157 same
11:42 3158 same
11:53 3159 same
12:05 3160 same
12:17 3161 same
J-P starts changing Target temperature at beginning of 3162
12:28 3162 same
12:40 3163 same
12:50 3164 same
J-P finished at beginning of 3164
12:59 3165 same
13:11 3166 same
13:21 3167 same
13:31 3168 same
13:41 3169 same
13:56 3170 same
14:05 ROC5 crash
14:08 3171 Production target_type=Loop 1 LH2 6 cm
14:10 20678 Production target_type=Loop 1 LH2 6 cm
14:19 3172 same
14:30 3173 same
14:43 3174 test run T3 with outer S1's off
14:47 3175 Production target_type=Loop 1 LH2 6 cm
14:58 3176 same
15:09 3177 same
15:20 3178 same
15:31 3179 same
15:42 3180 same
15:51 20679 same
5:52 3181 same